# Memory Gym: Partially Observable Challenges to Memory-Based Agents in Endless Episodes

### v1

@@ Abstract
Memory Gym introduces a unique benchmark designed to test Deep Reinforcement Learning agents, specifically comparing Gated Recurrent Unit (GRU) against Transformer-XL (TrXL), on their ability to memorize long sequences, withstand noise, and generalize. It features partially observable 2D environments with discrete controls, namely Mortar Mayhem, Mystery Path, and Searing Spotlights. These originally finite environments are extrapolated to novel endless tasks that act as an automatic curriculum, drawing inspiration from the car game “I packed my bag”. These endless tasks are not only beneficial for evaluating efficiency but also intriguingly valuable for assessing the effectiveness of approaches in memory-based agents. Given the scarcity of publicly available memory baselines, we contribute an implementation driven by TrXL and Proximal Policy Optimization. This implementation leverages TrXL as episodic memory using a sliding window approach. In our experiments on the finite environments, TrXL demonstrates superior sample efficiency in Mystery Path and outperforms in Mortar Mayhem. However, GRU is more efficient on Searing Spotlights. Most notably, in all endless tasks, GRU makes a remarkable resurgence, consistently outperforming TrXL by significant margins.
@@
Keywords: deep reinforcement learning, proximal policy optimization, memory, transformer-xl, gated recurrent unit
@@ 1 Introduction
In the evolving landscape of Deep Reinforcement Learning (DRL), the significance of memory-based tasks is increasingly evident. Our research introduces “Memory Gym”, a benchmark that emphasizes assessing memory capabilities in endless tasks, drawing inspiration from the classic car game “I packed my bag”. Central to our contribution is a Transformer-XL (Vaswani et al., 2017) driven baseline implementation for Proximal Policy Optimization (PPO) (Schulman et al., 2017), making it readily accessible to the community. This introduction delves into the intricate role of memory in decision-making agents, setting the stage for our key contributions. We conclude this section by outlining the structure and flow of the subsequent content in this paper.
1.1 “I Packed My Bag”: A Game to Explore Memory Limits
@@
Imagine embarking on a long-awaited family vacation, with the open road stretching ahead. As the car hums along, a lively atmosphere envelops the passengers, young and old alike, as they engage in a playful game of “I packed my bag”. However, as the game unfolds and the list lengthens, the passengers encounter mounting challenges illustrating the finite nature of their individual memories. Despite best efforts, varying strategies, and different capabilities, the weight of the growing list eventually leads to resignation. Due to the game’s ever-growing or endless concept, it serves as a revealing test of memory effectiveness, showcasing the limitations that individuals face when confronted with the increasing demands of retention and recall.
1.2 Memory in the Context of Decision-Making Agents
@@
Memory is not just a game – it is a critical tool for intelligent decision-making under imperfect information and uncertainty. Without the ability to recall past experiences, reasoning, creativity, planning, and learning may become elusive. In the realm of autonomously learning decision-making agents, the agent’s memory involves maintaining a representation of previous observations, a knowledge bank that grounds its next decision. Memory mechanisms, be it through recurrent neural networks (Rumelhart et al., 1986) or transformers (Vaswani et al., 2017), have enabled these agents to master tasks both virtual and real. For instance, DRL methods have conquered complex video games as Capture the flag (Jaderberg et al., 2018), StarCraft \@slowromancapii@ (Vinyals et al., 2019), and DotA 2 (Berner et al., 2019). Their success extends beyond virtual environments to real-world challenges as dexterous in-hand manipulation (Andrychowicz et al., 2020) and controlling tokamak plasmas (Degrave et al., 2022). The remarkable achievements of DRL agents empowered by memory often come coupled with substantial computational demands and additional techniques that extend beyond the scope of assessing an agent’s memory interaction abilities.
1.3 Contribution: Endless Environment Memory Benchmark
@@
In our prior research, we introduced Memory Gym (Pleines et al., 2023), an open-source benchmark, designed to challenge memory-based DRL agents to memorize events across long sequences, generalize, be robust to noise, and be sample efficient. Memory Gym encompasses three unique environments: Mortar Mayhem, Mystery Path, and Searing Spotlights. Each environment presents visual observations, discrete action spaces, and crucially, they cannot be mastered without memory, demanding frequent memory interactions.
Building on this foundation, we further enhance Memory Gym’s offerings by introducing endless tasks, drawing inspiration from the “I packed my bag” game. Such endless tasks, behaving as an automatic curriculum, are pivotal for not just evaluating efficiency but, more critically, assessing the effectiveness of memory mechanisms. As mirrored in the aforementioned game, capitulating to the ever-expanding list becomes inevitable. To our current understanding, no existing Deep Learning memory benchmark possesses such endless tasks targeting sequence models. Additionally, DRL tasks exhibit dynamic behaviors due to the evolving nature of the agent’s policy during training. As a result, episodes can yield vastly different outcomes. For example, episode lengths can vary significantly based on the decisions made, particularly when dealing with stochastic agents or environments.
It is important to clarify that “endless” should not be confused with “open-ended”. Memory Gym’s endless environments are not open-ended in nature; they possess clear objectives and defined structures. In contrast, open-ended environments offer agents freedom to develop unique skills and strategies without stringent constraints as demonstrated by the Paired Open-Ended Trailblazer algorithm (Wang et al., 2020) or platforms like Minecraft (Fan et al., 2022).
1.4 Contribution: Transformer-XL PPO Baseline
Many significant publications in memory-based DRL, particularly those utilizing transformers (Fortunato et al., 2019; Parisotto et al., 2020; Hill et al., 2021; Lampinen et al., 2021; Parisotto and Salakhutdinov, 2021), withhold their source code, impeding the replication of their findings and consequently hampering research progress. The absence of available baseline implementations is further discussed in section 5.2. In response to this challenge, we offer an accessible and easy-to-follow baseline implementation based on the PPO DRL algorithm (Schulman et al., 2017) and Transformer-XL (TrXL) (Vaswani et al., 2017). To allow for memory, we utilize TrXL as an episodic memory leveraging a sliding window approach.
@@ 1.4 Contribution: Transformer-XL PPO Baseline
In our experimental analysis, we benchmark both the TrXL and a recurrent agent based on the Gated Recurrent Unit (GRU) (Cho et al., 2014) across the original (finite) and the new endless environments of Memory Gym. Our findings indicate that TrXL performs best in terms of sample efficiency on Mystery Path and effectiveness on Mortar Mayhem. However, in the Searing Spotlights environment, GRU demonstrates better sample efficiency. To our greatest surprise, in the endless environments, GRU consistently surpasses TrXL in effectiveness by large margins.
@@ 1.5 Overview
This paper proceeds by delving into the properties and dynamics of the Memory Gym environments, highlighting Mortar Mayhem, Mystery Path, and Searing Spotlights. Additionally, we introduce their innovative endless counterparts, offering a fresh perspective on memory challenges. Subsequently, we outline the memory baselines that are tested against Memory Gym. Delving deeper, we describe the actor-critic model architecture, shedding light on the associated loss functions and the foundational components of both the GRU and TrXL baselines. Our experimental analysis follows, presenting empirical results from both finite and endless tasks within Memory Gym. Moreover, we discuss the intriguing aspect of TrXL’s suboptimal performance in endless tasks, offering insights, further experimentation, and future directions. Prior to our concluding remarks, we offer a comparative perspective by discussing related baselines and benchmarks in the memory-based DRL domain.
2 Memory Gym Environments
(a) Multi-Discrete
(b) Move Forward
(c) Rotate Left
(d) Rotate Right
Figure 1: The skin-colored agent in Memory Gym utilizes a multi-discrete action space (a) with distinct dimensions for horizontal (red arrows) and vertical (green arrows) velocity. Both dimensions include a no-op action, enabling movement in eight distinct cardinal directions (all arrows). The agent maintains a fixed speed of 3 pixels per step, with its forward direction determined by the combined horizontal and vertical velocity. The agent’s forward direction is visually represented by white-colored hands. Additionally, Mortar Mayhem and Mystery Path offer a grid-like locomotion system with a discrete action space of four actions: moving forward one tile at a time (b), rotating by 90 degrees left (c) or right (d), and a no-op action.
Mini-games of the commercial video game Pummel Party [^1] http://rebuiltgames.com/ inspired Memory Gym’s environments. The agent perceives all environments from a top-down perspective using $84\times 84$ RGB pixels, while its action space is multi-discrete, featuring two dimensions of size three as shown in Figure 1(a). Simpler grid-based locomotion variants are also incorporated in Mortar Mayhem and Mystery Path, as depicted in Figures 1(b) to 1(d). Subsequently, we dive into an in-depth exploration of the environments’ dynamics, highlighting their peculiarities towards memory, and above all, showcasing their distinctive endless design. Annex A presents the reset parameters of the enironments and their default values.
2.1 Mortar Mayhem
[…]
@@ 2.1 Mortar Mayhem
(c) Visual Feedback
Figure 2: Mortar Mayhem’s relevant entities are depicted in the annotated ground truth of the environment (a). The ground truth includes a green circle that designates the next target tile to move to. At the beginning of an episode, the commands are sequentially rendered onto the agent’s observation (b) while the agent remains immobile. Once all commands are displayed, the agent must navigate to the target tile within a specified time frame. Upon completion of the current command’s execution time, the agent’s success is evaluated, as illustrated in figure (c). This visual feedback is perceptible to the agent. Standing on red colored tiles means failure. Following a brief delay of a few steps during evaluation, the agent can proceed to the next command unless the episode has terminated due to failure or completion of the entire command sequence.
Mortar Mayhem (MM; Figure 2) takes place inside a grid-like arena and consists of two memory-dependent tasks. Initially, the agent is immobile and must memorize a sequence of ten commands (Clue Task), followed by executing each command in the observed order (Act Task). A command instructs the agent to move to an adjacent floor tile or remain at the current one. Failure to execute a command results in episode termination, whereas each successful execution yields a reward of $+0.1$. MM can be simplified to solely provide the Act Task (MMAct), where the command sequence is fully observable as a one-hot encoded feature vector. The Act Task necessitates frequent memory utilization since it must discern which commands have been executed to fulfill the subsequent ones, while considering a maximum of nine target tiles. To overcome this challenge, the agent can learn to track time (e.g. count steps) where a short-term memory seems sufficient.
MM’s difficulty can be scaled by several parameters, such as the number of commands, the show duration of a single command, and the delay between showing commands. Furthermore, the time allocated for executing a single command can be modified. These options can also be sampled upon resetting an episode. To simplify the difficulty further, the agent can be equipped with grid-like movement (MMGrid, MMActGrid), allowing it to move one tile at a time. Doing so, the agent is not commanded to move to diagonal tiles.
Episode lengths in MM vary based on the agent’s proficiency, lasting longer as the agent improves until reaching an upper bound (i.e. max episode length). The following equations determine the maximum episode length in MM by considering the durations and delays of the Clue Task and Act Task.
Clue Task $=(\text{Show Duration}+\text{Show Delay})\times\text{Command Count}$ (1)
Execution Length $=(\text{Execution Duration}+\text{Execution Delay})\times\text{Command Count}$ (2)
Act Task $=\text{Execution Length}-\text{Execution Delay}+1$ (3)
[…]
2.2 Endless Mortar Mayhem
(a) Observing and Executing Commands
(b) Agent Observation
Figure 3: In Endless Mortar Mayhem, the visualization and execution of commands are alternated (a) following the idea of an automatic curriculum. Commands (blue) are visualized only once in-between executing a new command and the previous ones. The arena (b) seamlessly occupies the entire screen, enabling the agent to reenter from the opposite side if it walks off.
Endless Mortar Mayhem (EMM; Figure 3) introduces a unique concept inspired by the toddler’s game ”I packed my bag…”. It extends the to-be-executed command sequence continuously, enabling potentially infinite sequences. Two significant modifications are made to MM to accommodate this feature. In the previous setup, the agent had to observe the entire sequence before executing it. However, in EMM, we adopt a different approach by alternating between command visualization and execution.
As shown in Figure 3(a), the agent only perceives one command during visualization. During execution, the agent must execute all previous commands along with the new one. An episode begins by displaying one command, followed by its execution. Subsequently, the next visualization reveals only the second command, requiring the agent to execute both the first and second commands. This automatic curriculum allows the to-be-executed command sequence to continually expand, while each new command is presented only once. A reward of $+0.1$ is signaled for every single successful command.
The second modification concerns the arena’s boundaries. In the original MM, the agent’s position would need to be reset after command execution due to certain commands being unavailable when the agent is adjacent to the arena’s boundaries. In EMM, we eliminate the borders and seamlessly scale the arena to the screen’s dimensions. Additionally, we implement a screen wrap feature that transforms the arena into a torus. Consequently, if the agent walks off the screen, it reappears on the opposite side without requiring relocation. Finally, the agent then executes the sequence of commands again from its new position.
2.3 Mystery Path
[…]
(a) Annotated Ground Truth
(b) Agent Observation
(c) Visual Feedback
Figure 4: The annotated ground truth depicts the relevant entities in Mystery Path (a). The walkable path is distinguished by its blue (origin), green (goal), and white color (path). Areas outside the path are designated as pits, represented by black and red colors, causing the agent to fall off and be relocated to the path’s origin on the next step. The initial randomly chosen red tiles ($A^{\ast}$ Walls) are only considered during path generation. The agent’s observation includes just itself (b). If the agent deviates from the path, a red cross provides visual feedback (c).
Mystery Path (MP; Figure 4) challenges the agent to traverse an invisible path within a grid-like level of dimension $7\times 7$. If the agent deviates from the path and thus falls down the pit, the agent is relocated to its origin during the next step. The episode terminates when the agent reaches the goal or exhausts its time budget of 512 steps. Reaching the goal yields a reward of $+1$. If a dense reward function is desired, the agent can be rewarded by $0.1$ for visiting previously unexplored tiles. To effectively navigate the environment, the agent must remember its steps on the path as well as the locations where it fell off. The invisible path is procedurally generated using the $A^{\ast}$ path-finding algorithm (Hart et al., 1968). Initially, the path’s origin is randomly sampled from the grid’s border, while the goal is placed on the opposing side in a random position. Instead of aiming for the shortest path, the cost function of $A^{\ast}$ is sampled to generate a diverse range of paths. The path generation process randomizes some cells as untraceable ($A^{\ast}$ walls), which adds variation to the generated paths. It is important to note that these walls are considered pits rather than physical obstacles.
MP can be simplified into a grid variant called MPGrid, which features a single discrete action space of four actions. In this variant, the agent’s time limit is reduced to 128 steps. Furthermore, MP can be made easier by including the path’s goal and origin in the agent’s observation. On the other hand, increasing the difficulty can be achieved by reducing the agent’s scale, decreasing its speed, or enlarging the path. In contrast to MM, episodes in MPGrid terminate earlier as the agent improves.
2.4 Endless Mystery Path
(a) Endless Path
(b) Agent Observation
Figure 5: In Endless Mystery Path, the path (a) is established by concatenating path segments, following a similar generation process as in MP. To avoid multiple neighboring path tiles, we add a transition tile (blue) between segments. Due to the constraints of representing an endless path within the observation space (b), we opt to visually fix the agent’s horizontal position. By rendering the path tiles behind the agent, the agent gains a sense of its horizontal motion, allowing it to perceive its progress along the path. Its horizontal position is adjustable, potentially obscuring more of the past path to increase the challenge.
Endless Mystery Path (EMP; Figure 5) introduces a never-ending path for the agent to traverse. Based on the original MP environment, we make several modifications to the path generation, the agent’s observation, and the terminal conditions.
The path generation in EMP (Figure 5(a)) involves concatenating path segments. Each segment is generated similar to MP, with the only difference being that the origin of a segment is based on the final tile of the previous segment. To ensure that path tiles do not have multiple neighbors, an additional tile is added between segments. The path is always generated from left to right. Consequently, we eliminate the agent’s ability to move left.
As the endless path cannot be fully captured in an $84\times 84$ RGB pixel observation, we visually fix the agent’s horizontal position. If nothing but the agent is shown in the observation, the agent lacks information about its horizontal motion. To address this, we render the path tiles behind the agent (Figure 5(a)), providing visual cues of the local horizontal position. For added difficulty, the visual horizontal position can be shifted further to the left, obscuring more information about the actual path.
Regarding terminal conditions, we aim to shorten episodes if the agent performs poorly. To achieve this, we implement three terminal conditions. Firstly, the agent is given a time budget of 20 steps to reach the next tile. Moving in the wrong direction or taking too long to progress within the budget leads to episode termination. The time budget is reset when the agent moves to the next tile or falls off and is relocated to the path’s origin. The second condition terminates the episode if the agent falls off before reaching its best progress. Lastly, the episode is terminated if the agent falls off at the same location for a second time. With these conditions in place, an episode can potentially last indefinitely if the agent consistently succeeds.
2.5 Searing Spotlights
[…]
(a) Annotated Ground Truth
(b) Agent Observation
(c) Visual Feedback
@@ 2.5 Searing Spotlights
Figure 6: The environment’s annotated ground truth (a) reveals information about all relevant entities in Searing Spotlights. The top rows of pixels display the agent’s remaining health points, its last action, and indicate whether a positive reward was received during the last step. The last action is encoded using two chunks and three colors, while the last positive reward is represented by two colors. Yellow circles indicate coins, while the somewhat rounded gray shape represents the exit. When all coins are collected, the exit turns green. The floor of the environment is checkered blue and white. In the agent’s observation (b), the spotlights play a role in revealing or hiding other entities. As an additional form of visual feedback (c), the blue floor tiles turn red if a spotlight detects the presence of the agent.
Searing Spotlights (Figure 6) presents a pitch-black environment to the agent, where information is revealed only through roaming and threatening spotlights. The agent starts with a limited number of health points (e.g. five) and loses one health point per step if hit by a spotlight. The episode terminates if the agent runs out of health points or exceeds the time limit of 256 steps. To evade the imminent threats, the agent must strategically hide in the darkness. In order to escape from the encroaching spotlights, the agent needs to remember its past actions and at least one previous position to deduce its current location. This requires the agent to carefully manage its health point budget to ascertain its location, ideally with minimal exposure. To avoid monotonous behavior, two additional tasks are introduced in the environment, requiring the agent to actively traverse the environment. The first task asks the agent to collect a predetermined number of coins, while the second task involves reaching an exit to finish the episode once all the coins have been collected. Both the exit and the coins are randomly positioned during the environment’s reset phase. Collecting a coin rewards the agent with $+0.25$, while successfully utilizing the exit grants a reward of $+1$. By utilizing its memory effectively, the agent shall be able to deduce its current location and recall the positions of the coins and the exit.
A simplified grid version of Searing Spotlights is not available. As a default simplification, the episode starts with perfect information, and after a few steps, the global light is gradually dimmed until off. Just like in MP, more successful episodes terminate earlier.
2.6 Endless Searing Spotlights
Based on just a few modifications, we extrapolate the concept of Searing Spotlights to Endless Searing Spotlights (ESS). We eliminate the exit task and alter the coin collection task by spawning a new coin whenever the agent collects one. The newly spawned coin is always visible for six frames and the agent has a time budget of 160 steps to collect the new coin. Lastly, we raise the agent’s health points from five to ten.
3 Memory Baselines

### v3

# Memory Gym: Towards Endless Tasks to Benchmark Memory Capabilities of Agents

@@ Abstract
Memory Gym presents a suite of 2D partially observable environments, namely Mortar Mayhem, Mystery Path, and Searing Spotlights, designed to benchmark memory capabilities in decision-making agents. These environments, originally with finite tasks, are expanded into innovative, endless formats, mirroring the escalating challenges of cumulative memory games such as “I packed my bag”. This progression in task design shifts the focus from merely assessing sample efficiency to also probing the levels of memory effectiveness in dynamic, prolonged scenarios. To address the gap in available memory-based Deep Reinforcement Learning baselines, we introduce an implementation that integrates Transformer-XL (TrXL) with Proximal Policy Optimization. This approach utilizes TrXL as a form of episodic memory, employing a sliding window technique. Our comparative study between the Gated Recurrent Unit (GRU) and TrXL reveals varied performances across different settings. TrXL, on the finite environments, demonstrates superior sample efficiency in Mystery Path and outperforms in Mortar Mayhem. However, GRU is more efficient on Searing Spotlights. Most notably, in all endless tasks, GRU makes a remarkable resurgence, consistently outperforming TrXL by significant margins.
@@
Keywords: deep reinforcement learning, proximal policy optimization, memory, transformer, recurrence
@@ 1 Introduction
Imagine embarking on a long-awaited family vacation, with the open road stretching ahead. As the car hums along, a lively atmosphere envelops the passengers, young and old alike, as they engage in a playful game of “I packed my bag”. Each family member takes turns adding an item to an imaginary bag, attempting to remember and recite the growing list correctly. However, as the game unfolds and the list lengthens, the passengers encounter mounting challenges illustrating the finite nature of their individual memories. Despite best efforts, varying strategies, and different capabilities, the weight of the growing list eventually leads to resignation.
@@
Memory is not just a game – it is a critical tool for intelligent decision-making under imperfect information and uncertainty. Without the ability to recall past experiences, reasoning, creativity, planning, and learning may become elusive. In the realm of autonomously learning decision-making agents as Deep Reinforcement Learning (DRL), the agent’s memory involves maintaining a representation of previous observations, a knowledge bank that grounds its next decision. Memory mechanisms, be it through recurrent neural networks (Rumelhart et al., 1986) or transformers (Vaswani et al., 2017), have enabled these agents to master tasks both virtual and real. For instance, DRL methods have conquered complex video games as Capture the flag (Jaderberg et al., 2018), StarCraft \@slowromancapii@ (Vinyals et al., 2019), and DotA 2 (Berner et al., 2019). Their success extends beyond virtual environments to real-world challenges as dexterous in-hand manipulation (Andrychowicz et al., 2020) and controlling tokamak plasmas (Degrave et al., 2022).
1.1 Endless Tasks Benchmark Memory Effectiveness and not just Efficiency
Cumulative memory games, such as “I packed my bag”, introduce an automatic curriculum through their inherently progressive and unbounded nature. The game’s expansion depends on the participants’ continuous success. Exploiting the concept of ‘endlessness’ in tasks can provide deeper insights into the effectiveness of decision-makers when facing escalating demands of memory retention and recall. Such challenges inevitably test the limits of an agent’s memory. Here, ‘effectiveness’ refers to the ability to attain higher scores, indicative of more powerful memory capabilities. Contrasting this, traditional DRL memory benchmarks are centered around sample efficiency, influenced by the defined boundaries and fixed goals of finite tasks, which do not accommodate perpetual tasks.
1.2 Contributions: Novel Memory Benchmark and Transformer-XL Baseline
@@
In our prior research, we introduced Memory Gym (Pleines et al., 2023), an open-source benchmark, designed to challenge memory-based DRL agents to memorize events across long sequences, generalize, be robust to noise, and be sample efficient. Memory Gym encompasses three distinct finite environments: Mortar Mayhem, Mystery Path, and Searing Spotlights. Each environment presents visual observations, multi-discrete action spaces, and crucially, they cannot be mastered without memory, demanding frequent memory interactions. Building on this foundation, we advance Memory Gym by evolving each environment into an endless format, inspired by the “I packed my bag” game, to further test and develop memory capabilities in DRL agents. To our current understanding, no existing DRL memory benchmark possesses such endless tasks.
We further contribute an open-source and easy-to-follow baseline implementation of Transformer-XL (TrXL) (Dai et al., 2019) based on the widely-used DRL algorithm Proximal Policy Optimization (PPO) (Schulman et al., 2017). A recurring theme in the realm of memory-based DRL is the lack of accessible and reproducible baselines. Many significant works, particularly those utilizing transformers (Fortunato et al., 2019; Parisotto et al., 2020; Hill et al., 2021; Lampinen et al., 2021; Parisotto and Salakhutdinov, 2021), have not made their implementations publicly available. As implementation details of DRL algorithms are vital (Engstrom et al., 2020; Andrychowicz et al., 2021; Huang et al., 2022b), this lack of transparency poses significant challenges for the research community, hindering progress and reproducibility. Therefore, our accessible implementation can serve as a beacon for the community fostering further research.
@@
In our experimental analysis, we benchmark a TrXL and a recurrent agent, based on the Gated Recurrent Unit (GRU) (Cho et al., 2014), across the original (finite) and the new endless environments of Memory Gym. Our findings indicate that TrXL performs best in terms of sample efficiency on Mystery Path and effectiveness on Mortar Mayhem. However, in the Searing Spotlights environment, GRU demonstrates better sample efficiency. To our greatest surprise, in the endless environments, GRU consistently surpasses TrXL in effectiveness by large margins.
@@ 1.3 Overview
This paper begins with related work, followed by illustrating the dynamics in Memory Gym’s environments, particularly focusing on their endless formats. We detail the actor-critic model architecture, including the loss functions and key elements of the TrXL baseline. The subsequent section presents our experimental analysis, showcasing results from both finite and endless tasks. We further explore TrXL’s underperformance in endless tasks, providing insights and suggestions for future research. Before concluding, we falsify our initial hypothesis about recurrence’s vulnerability to spotlight perturbations in Searing Spotlights, emphasizing the significant impact of not normalizing estimated advantages during agent optimization.
2 Related Work
Previous studies have explored the use of memory-based agents in a variety of tasks. Some of them are originally fully observable but are turned into partially observable Markov Decision Processes by adding noise or masking out information from the agent’s observation space (Hausknecht and Stone, 2015; Heess et al., 2015; Meng et al., 2021; Shang et al., 2021). Subsequently we give a coarse overview of recently used benchmarks chronologically.
Deepmind Lab 30 (Beattie et al., 2016) presents a collection of 30 procedurally generated first-person 3D environments. Starting from a general motor-control navigation task and visual observations, each environment poses a different goal such as collecting fruit, playing laser tag or traversing a maze. Pašukonis et al. (2023) have identified that agents can exploit the environments’ skyboxes, which reduces the demands on memory.
PopGym (Morad et al., 2023), published at ICLR 2023 alongside Memory Gym (Pleines et al., 2023), is a benchmark with 15 environments classified as diagnostic, control, noisy, game, or navigation. These are designed for rapid convergence, offering vector instead of visual observations.
Memory Maze (Pašukonis et al., 2023) showcases procedurally generated 3D mazes, where agents, from a first-person perspective, are tasked with locating specific objects as requested by the environment. Additionally, the authors have collected a dataset tailored for offline reinforcement learning.
Memory Gym’s environments offer a unique feature that sets them apart from the aforementioned benchmarks: an endless design tailored to rigorously test memory capabilities. It is important to clarify that ‘endless’ should not be confused with ‘open-ended’. Memory Gym’s endless environments are not open-ended in nature; they possess clear objectives and defined structures. In contrast, open-ended environments offer agents freedom to develop unique skills and strategies without stringent constraints as demonstrated by the Paired Open-Ended Trailblazer algorithm (Wang et al., 2020) or platforms like Minecraft (Fan et al., 2022).
3 Endless Memory Tasks Inspired by “I Packed My Bag”
Prior to the underlying work, we introduced Memory Gym, a set of finite tasks benchmarking agents to strongly depend on frequent memory interactions (Pleines et al., 2023). Memory Gym consists of the environments Mortar Mayhem, Mystery Path, and Searing Spotlights. All of them expose visual observations, capturing an area of, $84\times 84$ RGB pixels, to the agent and allow for multi-discrete actions. Importantly, the original environments pose a single (or finite) goal. In this section, we delve into the intricacies of extrapolating the originally finite tasks to endless ones. Annex A presents the reset parameters to configure and scale the environments.
3.1 Endless Mortar Mayhem
(a) Agent Observation
(b) Command Observation and Execution
Figure 1: Endless Mortar Mayhem (a) is observed from a top-down view and its arena occupies the entire screen. If the agent, depicted as a circle with white paws, walks off, it re-enters from the opposite side. Episodes endlessly expand by alternately visualizing and executing commands (b), behaving as an automatic curriculum.
The finite variant of Mortar Mayhem takes place inside a grid-like arena and comprises two tasks. Initially, the agent is immobile and must memorize a sequence of ten commands, followed by executing each command in the observed order. A command instructs the agent to move to an adjacent floor tile or remain at the current one. Failure to execute a command results in episode termination, whereas each successful execution yields a reward of $+0.1$.
Endless Mortar Mayhem (Figure 1) extends the to-be-executed command sequence continuously, enabling potentially infinite sequences. To accommodate this, the command visualization and execution are alternated. As shown in Figure 1(b), an episode begins by displaying one command, followed by its execution. Subsequently, the next visualization reveals only the second command, requiring the agent to execute both the first and second commands. This automatic curriculum allows the to-be-executed command sequence to continually expand, while each new command is presented only once.
3.2 Endless Mystery Path
(a) Agent Observation
(b) Endless Path
Figure 2: In Endless Mystery Path (a), the agent observes only a segment of the environment at a time. By rendering the path tiles past it, the agent acquires a sense of horizontal movement, even though its horizontal position remains visually constant. Figure (b) showcases an example of a procedurally generated endless path.
The core concept of the original Mystery Path environment requires the agent to navigate an invisible path until reaching its end. If the agent strays off the path, it is relocated to the path’s origin. To succeed, the agent must consistently memorize and recall its steps along the path, as well as the locations where it fell off.
Endless Mystery Path (Figure 2) introduces a never-ending path, which is always generated from left to right. Consequently, we eliminate the agent’s ability to move left. As the endless path cannot be fully captured in an $84\times 84$ RGB pixel observation, we visually fix the agent’s horizontal position. If nothing but the agent is shown in the observation, the agent lacks information about its horizontal motion. To address this, we render the path tiles behind the agent, providing visual cues of the local horizontal position.
Regarding terminal conditions, we aim to shorten episodes if the agent performs poorly. Firstly, the agent is given a time budget of $20$ steps to reach the next tile. This budget is reset when the agent moves to the next tile or falls off. Another condition terminates the episode if the agent falls off before reaching its best progress. Lastly, the episode is terminated if the agent falls off at the same location for a second time. With these conditions in place, an episode can potentially last indefinitely if the agent consistently succeeds.
In terms of rewards, the agent earns $+0.1$ for each previously unvisited tile of the path it reaches. While this may initially appear as a dense reward function, any misstep that causes the agent to fall off means it cannot reclaim those rewards. The agent must retrace its steps and regain its prior progress to continue accumulating rewards. The longer it takes to catch up, the sparser the rewards become.
3.3 Endless Searing Spotlights
Health BarLast ActionLast RewardAgentSpotlightCoinRefer to caption
@@ 3.3 Endless Searing Spotlights
Figure 3: In Endless Searing Spotlights’ annotated ground truth, the top rows of pixels display the agent’s remaining health points, its last action, and indicate whether a positive reward was received during the last step. The last action is encoded using two chunks and three colors, while the last positive reward is represented by two colors. In the agent’s observation (b), the spotlights play a role in revealing or hiding other entities.
Endless Searing Spotlights (Figure 3) features a pitch-black environment illuminated only by moving and threatening spotlights. The agent starts with ten health points, losing one each time it is hit by a spotlight. At the beginning of each episode, the environment is fully lit, but it quickly dims to darkness within a few steps. The episode ends if the agent’s health points are depleted. To survive, the agent must cleverly use the darkness as cover, relying on memory to remember past actions and previous positions to infer its current location. The agent’s observation space includes its last action, current health, and any positive reward received in the previous step.
To mitigate monotonous behaviors, a coin collection task is incorporated. In the finite version, the agent aims to collect a randomly placed coin, worth $+0.25$, and then exit the level for an additional $+1$ reward. Effective memory use enables the agent to track its location, coin positions, and the exit. In the endless concept, the exit is removed, and the coin collection mechanism is revamped to spawn a new coin each time one is collected. This coin remains visible for six frames, while the agent has 160 steps to collect it. Distinct from the finite version, spotlights now appear at a steady rate, rather than intensifying over time.
4 Transformer-XL Baseline